### v1

# Looking into the Jet Cone of the Neutrino-Associated Very High Energy Blazar PKS 1424+240

Abstract
The acceleration process of massive particles as well as the production of very high energy (VHE) photons and neutrinos remains a fundamental challenge in astrophysics. We investigate the parsec-scale properties of the blazar PKS 1424+240, that has been selected on the basis of both gamma-ray and neutrino VHE radiation. We analyze VLBA observations of this BL Lac object, stacking 42 polarization-sensitive images collected in 2009–2025 to enhance the signal and reveal persistent parsec-scale structure. Our observations indicate that this object is viewed inside the jet cone, very close to the axis of its relativistic jet, with a viewing angle of $<0.6^{\circ}$. This effectively maximizes Doppler boosting to values $\sim 30$ and enhances both electromagnetic and neutrino emission in the observer’s direction. Based on polarimetric observations, we unambiguously detect a net toroidal component of the jet’s magnetic field. Blazars with very small jet viewing angles offer a solution to the longstanding mismatch between Doppler factors inferred from VLBI and those derived from VHE observations — the so-called ‘Doppler factor crisis’. We show that relativistic beaming plays the critical role in the gamma-ray and neutrino emission of blazars, with direct implications for models of their multi-messenger emission.
1 Introduction
Understanding how the universe produces photons and neutrinos at the highest observable energies is of broad interest, as it directly relates to fundamental physics that can only be explored through astronomy. Active galactic nuclei (AGN) are strong emitters of very high energy (VHE) photons (e.g., J. Albert et al., 2007; F. Aharonian et al., 2007) and have been proposed as efficient proton accelerators (starting with V. Berezinsky, 1977), capable of producing neutrinos at TeV–PeV and beyond. Their observed properties are strongly influenced by relativistic beaming effects that also introduce significant observational biases. For example, flux-density-limited samples of radio Very Long Baseline Interferometry (VLBI) selected AGN are dominated by distant beamed blazars with a median redshift of $z\sim 1$ (L. Y. Petrov & Y. Y. Kovalev, 2025). However, the role of beaming at very high energies ($\gtrsim 100$ GeV) is not firmly established. Indeed, while fast variability of VHE gamma-rays implies high values of the Doppler factor $\delta$, the lack of strong apparent superluminal motions in VLBI observed jets suggests the opposite. This is known as the ‘Doppler factor crisis’ (see, e.g., J. Albert et al., 2007; F. Aharonian et al., 2007; B. G. Piner & P. G. Edwards, 2018). Breakthroughs in this area have been rare over the past 35 years, with the Doppler factor crisis posing a significant barrier to progress. The neutrino emission from distant blazars is expected to be beamed, with proton acceleration occurring preferentially along the jet direction (e.g., IceCube Collaboration et al., 2018; A. V. Plavin et al., 2025). In contrast, weaker AGN may emit neutrinos (R. Abbasi et al., 2022) without any preferred direction.
Figure 1: The IceCube skymap cutout surrounding the radio position of the PKS 1424+240 blazar. The color scale represents the local p-value from the 9-year maximum likelihood analysis performed by R. Abbasi et al. (2022).
With a redshift $z=0.605$ (S. Paiano et al., 2017), the BL Lac object PKS 1424+240 (common name OQ 240 (catalog )) is the most distant blazar detected in quiescent VHE (e.g., J. Aleksić et al., 2014; P. Padovani et al., 2022). The all-sky analysis performed by R. Abbasi et al. (2022) has identified it as the possible source of the second-highest neutrino excess in the predefined source list in the Northern sky (see Figure 1). The best-fit neutrino spectrum for PKS 1424+240 is soft with a spectral index of $\gamma=3.5$ (R. Abbasi et al., 2022). The IceCube signal is dominated by neutrinos with relatively low, $\lesssim 10$ TeV, energies. The time-integrated IceCube analysis is predominantly sensitive to persistent emission. Together with off-flare VHE detections in gamma-rays, this motivates the study of time-independent properties of PKS 1424+240. Here, we address the question, which general property of PKS 1424+240 may be responsible for its exceptional VHE photon and neutrino fluxes.
We use radio VLBI, which is the only technique to date that is able to directly probe the sites of neutrino and gamma-ray production, resolving structures within several parsecs from the central engine. For the redshift of PKS 1424+240, the projected angular scale of 1 mas corresponds to 6.70 pc, assuming the E. Komatsu et al. (2009) cosmological parameters. A quick examination of PKS 1424+240 has revealed that its jet has an exceptionally large apparent opening angle $\varphi_{\mathrm{app}}$. A. B. Pushkarev et al. (2017) found $\varphi_{\mathrm{app}}=65\arcdeg$, among the most extreme values observed for blazars. Geometrically, this indicates a small viewing angle $\theta$ to the line of sight and, consequently, strong Doppler boosting. To further investigate the source, we updated its stacked 15 GHz image by the Very Long Baseline Array (VLBA) using all available data and increasing the number of stacked epochs by about ×$\times 4$. The results of our serendipitous findings are presented in this paper.
2 VLBA polarization imaging
2.1 Long-term 15 GHz MOJAVE observations: image stacking
Stokes I and polarization observation of PKS 1424+240 in the MOJAVE program [^1] https://www.cv.nrao.edu/MOJAVE/ (MOJAVE: Monitoring Of Jets in Active galactic nuclei with VLBA Experiments, see M. L. Lister et al., 2018, and references therein) began in May 2009 and we report results through January 2025. To increase the dynamic range of a restored map, a set of single-epoch images is combined and averaged, i.e., stacked, by aligning them using the VLBI core position derived from visibility model fitting (e.g., M. L. Lister et al., 2021). Stacking not only improves the sensitivity and dynamic range but also effectively reconstructs more complete source morphology, particularly in the low-brightness areas. It can also reveal the whole jet channel, not just regions of enhanced radiation at a given epoch. It is beneficial for studies of jets in Stokes I and polarized emission (e.g. A. B. Pushkarev et al., 2017, 2023; Y. Y. Kovalev et al., 2020; D. I. Zobnina et al., 2023).
We made stacked maps of PKS 1424+240 in Stokes I, Q, and U (Figure 2), using 42 available epochs that span approximately 15 years, following a procedure described in A. B. Pushkarev et al. (2023). For deep CLEANing in Stokes I, Q, and U, we applied the algorithm with the residual entropy-based stopping criterion developed by D. C. Homan et al. (2024). The linear polarization image has been corrected for Ricean bias (J. F. C. Wardle & P. P. Kronberg, 1974). We also applied a debiasing procedure to the CLEAN images to effectively suppress the sidelobe artifacts from deconvolution (Appendix B in A. B. Pushkarev et al., 2023).
Figure 2: VLBA stacked image for the blazar PKS 1424+240 at 15 GHz. Stokes I is shown by contours, the first contour corresponds to 3×3$\times$3 × the image rms level. Linear polarization intensity is presented by color, directions of electric vector position angle (EVPA) — by sticks. The observation dates of the 42 images used in the stacked image are indicated in the inset. A circular restoring beam is shown at the full width at half maximum (FWHM, 0.8 mas) level in the bottom left corner. FITS files of the stacked Stokes I, Q, and U images are published as electronic supplementary material. We also supply an animation of Stokes I cumulative stacking.
This debiasing procedure has significantly reduced imaging artifacts, and we argue that the weak structures observed around the core of PKS 1424+240 in the stacked images (Figure 2) represent genuine reconstructed jet emission. The northern declination of PKS 1424+240 ensures robust VLBA $uv$-coverage with no significant gaps. The weak polarized emission stacks coherently over multiple epochs, and is not expected to result from residual polarization leakage effects. The reduced levels of detectable 15 GHz Stokes I emission in the North-West direction next to the bright core component possibly results from residual effects of self-calibration which only included emission to the South-East in the self-calibration model in most individual epochs.
In order to check for consistency of the stacked 15 GHz image, we have processed archival 1.6 GHz VLBA data for PKS 1424+240, experiment BS257A, epoch 9 May 2017. At ten times lower frequency, the VLBA is substantially more sensitive to distant optically thin jet emission and probes angular scales ten times larger than our MOJAVE stacked image. With a total on-source time of about 3 hr, these single epoch low-frequency observations reach a high dynamic range of $\sim 7,000:1$ and confirm the emission around the VLBI core, including the North-West direction (Figure 4). We note that our stacked image at 24 GHz also detects this emission (Figure 5).
2.2 Rotation Measure at 15, 24, and 43 GHz
Between August 2019 and June 2021, the MOJAVE monitoring conducted monthly 15 GHz, 24 GHz, and 43 GHz band polarimetric observations of 25 AGN (including PKS 1424+240), totaling 23 epochs, in order to measure the rotation measure (RM; full results will be presented by Livingston and the MOJAVE team, in prep.). In Figure 5 we show the mean RM of PKS 1424+240 across the two years of observations. This object is one of the lowest in total magnitude of observed RM for its core, with a median magnitude across all epochs of $302\pm 108\,\mathrm{rad\,m^{-2}}$, compared to a population median magnitude of $1026\pm 475\,\mathrm{rad\,m^{-2}}$. We also note that the median pixel-by-pixel standard deviation of RM for the core is $514\pm 57\,\mathrm{rad\,m^{-2}}$, making it larger than the median magnitude of RM for the core. Interestingly, across the two years, the core of PKS 1424+240 does not vary greatly in the magnitude of RM, sitting in the bottom $\mathrm{25^{th}}$ percentile of time variability of sources. However, the jet of PKS 1424+240 has a time variability of RM similar to other sources within the sample. We also detect frequency dependent depolarization, which is typical of most sources within our sample. This combined with the low magnitude RM suggests that the Faraday rotation that we do see is external to the jet.
As RM data is not available for all 42 epochs used in Figure 2, and individual epochs do not have RM data for the full jet region, we cannot apply a precise correction for the Faraday effect on EVPA. The median magnitude of RM for PKS 1424+240 corresponds to a EVPA rotation of $\sim 6^{\circ}$, which is comparable to our EVPA errors from calibration and imaging.
The low magnitude and variability of RM especially close to the core of PKS 1424+240 are indicative of either a weak magnetic field or low thermal electron density along the line-of-sight. We may be seeing the ‘sweeping away’ of the magneto-ionic medium, this would result in a reduction in the magnitude of RM. This effect would be more prominent within the center of the jet, while the edges would be less affected, which is supported by the magnitude and variability of RM being typical of other sources for the jet region of PKS 1424+240.
3 Looking into the jet cone
3.1 Observational signatures and magnetic field
Figure 3: The Eye of Sauron. Stacked VLBA linear polarization image with magnetic field direction incorporated by the Linear Integral Convolution. The circular restoring beam is shown at the FWHM level of 0.8 mas in the bottom left corner.
We have collected striking observational signatures of a jet that is viewed inside its cone. The stacked Stokes I and linear polarization images reveal a structure in the southeast direction, while significant emission is also present in all other directions around the core region in Stokes I and linear polarization (Figure 2). This is extended even to deprojected kiloparsec scales, see Figure 4. Furthermore, the stacked polarization image exhibits a remarkably uniform distribution of EVPAs; it presents a pattern of “diverging rays”, which start from the core. The low rotation measure and its minimal variation suggest that the jet is propagating toward the observer, reducing the density of the interstellar medium along its path. The low jet speeds, observed by VLBI (subsection 3.2), are consistent with a small jet viewing angle, $\theta$.
We reconstruct the magnetic field structure, implementing a $90^{\circ}$ EVPA rotation across the entire map. This assumes that RM as well as relativistic effects do not affect the observed projected magnetic field direction significantly, see subsection 2.2 and M. Lyutikov et al. (2005). The results are shown in Figure 3, which (after J. R. R. Tolkien) we call ‘the Eye of Sauron’ due to its striking nature. Linear Integral Convolution (LIC; B. Cabral & L. C. Leedom, 1993) is applied to the polarized intensity to visualize the morphology of magnetic field lines by locally blurring intensity variations along the magnetic field direction, effectively tracing its structure. The image reveals a significant toroidal magnetic field component of a current-carrying jet, which flows almost directly toward Earth. This is consistent with a widely discussed overall helical magnetic field structure (R. Blandford et al., 2019).
3.2 Basic jet parameters
M. L. Lister et al. (2021) have measured apparent kinematics of two robust components in the emission of PKS 1424+240. Their apparent speed, measured in units of the speed of light, is $\beta_{\mathrm{app,1}}=2.83\pm 0.89$ and $\beta_{\mathrm{app,2}}=1.91\pm 0.18$. These moderate values are consistent with other measurements for the same source (e.g., P. Padovani et al., 2022; E. Kun & A. Medveczky, 2023). A median intrinsic full opening angle of the jet was estimated by A. B. Pushkarev et al. (2017) to be $\alpha_{\mathrm{int}}=1.2^{\circ}$ for Fermi-detected AGN in MOJAVE sample. We assume this value for PKS 1424+240. Looking into the jet cone condition is defined by the requirement for the jet viewing angle $\theta<\alpha_{\mathrm{int}}/2$. This requires the Doppler factor to reach its maximum possible value of twice the Lorentz factor, $\delta\approx 2\Gamma$.
We apply a simplified assumption that the jet is observed at a half of its half opening angle, $\theta=\alpha_{\mathrm{int}}/4$. This is in agreement with simulations by Pashchenko et al. (in prep.). We also assume a relation between $\alpha_{\mathrm{int}}$ and $\Gamma$, as deduced observationally by A. B. Pushkarev et al. (2017) and predicted by hydrodynamical and magnetic acceleration models of relativistic jets (R. D. Blandford & A. Königl, 1979; S. S. Komissarov et al., 2007; E. Clausen-Brown et al., 2013). As a result, we find that the following set of parameters for a basic relativistic jet model are consistent with the above values and considerations: $\Gamma=16$ (plasma flow speed $\beta=0.998$), $\alpha_{\mathrm{int}}=1.2^{\circ}$, $\theta=0.3^{\circ}$, $\beta_{\mathrm{app}}=2.8$, and Doppler factor $\delta=32$. The Lorentz factor is within its typical values from MOJAVE population modeling (M. L. Lister et al., 2019), as expected from the chosen conservative parameter assumptions. The Doppler factor is three times higher than the median values for the MOJAVE sample (M. L. Lister et al., 2021; D. C. Homan et al., 2021), which agrees with recent finding of A. V. Plavin et al. (2025) for neutrino-selected blazars. Note that $\Gamma$ and $\delta$ estimates are not highly sensitive to the exact value of $\theta$, as long as the line of sight remains within the jet cone. Assuming a jet viewing angle in the range $0.1^{\circ}<\theta<1.0^{\circ}$ and using the measured $\beta_{\mathrm{app}}$ values, the Doppler factors lie within the range $1652>\delta>$ 16. Although the geometric probability of a near-zero viewing angle is low, the fraction of such jets in VLBI flux density-selected samples has been shown to be on the order of a few percent due to Doppler bias (M. L. Lister et al., 2019).
Due to strong projection effects and synchrotron opacity, the jet is expected to become visible many parsecs downstream from its true base. As a result, its observed radio core appears less active and variable than typically expected for highly boosted blazars. Moreover, the core and jet emission are superimposed, resulting in a significant underestimation of the intrinsic core brightness temperature, which in turn leads to an incorrect estimation of jet parameters in D. C. Homan et al. (2021).
4 Emission of high energy photons and neutrinos: relativistic beaming
[…]
We note that the explanation of the VHE gamma-ray flux within a leptohadronic model of M. Cerruti et al. (2017) requires the Doppler factor of $\sim 30$, consistent with our observations. The expected associated neutrino flux is of order of that inferred from IceCube observations. The neutrino production in the blazar radio core may proceed through the $p\gamma$ mechanism, adopted by A. V. Plavin et al. (2021); O. Kalashev et al. (2023) to explain TeV to PeV neutrino production. Energetic protons, required for the neutrino and, in the hadronic scenarios, gamma-ray production, may be accelerated either in the vicinity of the central black hole (K. Ptitsyna & A. Neronov, 2016) or at the border surface between the fast spine and the slower sheath in the jet (e.g., F. Tavecchio & G. Ghisellini, 2015). In any case, the specific geometry of the jet in PKS 1424+240, which is closely aligned to the line-of-sight, provides for an additional Doppler enhancement of its multi-wavelength and multi-messenger fluxes. This boosting makes the object persistently bright, maintaining a high average flux — placing it among the top 1 % of gamma-ray sources (J. Ballet et al., 2023), and the brightest blazar in terms of high-energy neutrino emission (R. Abbasi et al., 2022). Future observations of this and other VHE emitting blazars are needed to pave the way to a quantitative model of neutrino production in jets, and to better understand the role of beaming in the gamma-ray emission.
5 Summary
Supermassive black holes at the heart of distant active galaxies are believed to be among the most powerful particle accelerators in the Universe. They produce VHE photons and serve as possible sources of cosmic neutrinos. The mechanism responsible for accelerating massive particles to relativistic energies remains an open question in modern astrophysics, with powerful extragalactic jets being a key candidate. A longstanding discrepancy—known as the Doppler factor crisis—arises from the mismatch between low apparent plasma speeds observed with VLBI and the highly relativistic flows required to explain the emission. We use 16 years of 2 cm VLBI observations to uncover properties of the parsec-scale jet of the blazar PKS 1424+240, which is exceptional in its VHE gamma-ray and probable neutrino emissions. A significant toroidal magnetic-field is detected by a clear signature in linear polarization, revealing a current-carrying jet flowing almost directly toward us. We determine that the jet is observed within its cone, at a viewing angle below half a degree. This geometry maximizes relativistic boosting while minimizing apparent speed, thereby resolving the Doppler factor crisis and suggesting extreme relativistic beaming for this VHE photon and neutrino emitting blazar. Our findings for PKS 1424+240 pave the way towards better understanding of high-energy photon and neutrino emission from blazar jets.
Observing jets within their plasma cone provides a unique opportunity to determine their intrinsic parameters and reconstruct the projected magnetic field structure. Moreover, such jets may represent a distinct class of AGN with both enhanced VHE gamma-ray and neutrino emission. From Monte Carlo simulations of the complete, VLBI-selected MOJAVE 1.5JyQC sample, we estimate that only a few percent of the jets are viewed within a degree to our line of sight (M. L. Lister et al., 2019). The results of a full MOJAVE sample analysis will be presented separately by Pushkarev et al. (in prep.).
We thank Andrei Lobanov, Yannis Liodakis, and Tigran Arshakian for productive discussions and comments on the manuscript. This work is part of the MuSES project which has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No 101142396). The rotation measure analysis by JDL is supported by the M2FINDERS project, which has received funding from the ERC under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101018682). AVP is a postdoctoral fellow at the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (grants 60477, 61479, 62286) and the Gordon and Betty Moore Foundation (grant GBMF-8273). The work of SVT and ABP is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541. The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research made use of the data from the MOJAVE database maintained by the MOJAVE team (M. L. Lister et al., 2018).
Appendix A Supplementary Figures
Figure 4: 1.6 GHz VLBA image of PKS 1424+240, epoch 9 May 2017. A circular restoring beam is shown in the left bottom corner at the FWHM level of 8.17 mas. It supports the viewing angle being less than a half-opening angle even at large (kiloparsec, deprojected) spatial scales probed by the low-frequency observations. Stokes I image FITS file is published electronically as a paper supplementary material.
Figure 5: Stacked mean RM map of 15/24/43 GHz VLBA data for PKS 1424+240 in the frame of the observer across 23 epochs from 15 August 2019 to 25 June 2021 in color scale. The contours show the mean 24 GHz Stokes I; the first contour corresponds to 3×3$\times$3 × the image rms level, with contours increasing by a factor of 2. The grey circle indicates the FWHM level of the circular restoring beam. The map shows the magnitude of RM is low and similar across the source.

### v2

@@
 Abstract
-The acceleration process of massive particles as well as the production of very high energy (VHE) photons and neutrinos remains a fundamental challenge in astrophysics.
-We investigate the parsec-scale properties of the blazar PKS 1424+240, that has been selected on the basis of both gamma-ray and neutrino VHE radiation.
-We analyze VLBA observations of this BL Lac object, stacking 42 polarization-sensitive images collected in 2009–2025 to enhance the signal and reveal persistent parsec-scale structure.
-Our observations indicate that this object is viewed inside the jet cone, very close to the axis of its relativistic jet, with a viewing angle of $<0.6^{\circ}$.
+Context. The acceleration process of massive particles as well as the production of very high energy (VHE) photons and neutrinos remains a fundamental challenge in astrophysics.
+Aims.
+We investigate the parsec-scale jet structure and magnetic field of the blazar PKS 1424+240, that has been selected on the basis of strong VHE gamma-ray emission and identified with one of the highest peaks in the IceCube 9-year neutrino sky.
+Methods. We analyze 15 GHz VLBA observations of this BL Lac object, stacking 42 polarization-sensitive images collected in 2009–2025 to enhance the signal and reveal persistent parsec-scale structure.
+Results.
+Our observations uncover a rare scenario.
+The object is viewed inside the jet cone, very close to the axis of its relativistic jet, with a viewing angle of $<0.6^{\circ}$.
 This effectively maximizes Doppler boosting to values $\sim 30$ and enhances both electromagnetic and neutrino emission in the observer’s direction.
-Based on polarimetric observations, we unambiguously detect a net toroidal component of the jet’s magnetic field.
-Blazars with very small jet viewing angles offer a solution to the longstanding mismatch between Doppler factors inferred from VLBI and those derived from VHE observations — the so-called ‘Doppler factor crisis’.
+Based on polarimetric observations, we unambiguously detect a net toroidal component of the jet’s magnetic field, indicating a current carrying jet flowing almost directly towards our line of sight.
+Conclusions.
+Blazars with very small jet viewing angles offer a solution to the longstanding mismatch between Doppler factors inferred from low VLBI apparent jet speed and those derived from VHE observations — the so-called “Doppler factor crisis”.
 We show that relativistic beaming plays the critical role in the gamma-ray and neutrino emission of blazars, with direct implications for models of their multi-messenger emission.
+Key Words.:
+Neutrinos – Radio continuum: galaxies – Galaxies: jets – BL Lacertae objects: individual: PKS 1424+240
+[…]
 1 Introduction
 Understanding how the universe produces photons and neutrinos at the highest observable energies is of broad interest, as it directly relates to fundamental physics that can only be explored through astronomy.
-Active galactic nuclei (AGN) are strong emitters of very high energy (VHE) photons (e.g., J.
-Albert et al., 2007; F.
-Aharonian et al., 2007) and have been proposed as efficient proton accelerators (starting with V.
-Berezinsky, 1977), capable of producing neutrinos at TeV–PeV and beyond.
+Active galactic nuclei (AGN) are strong emitters of very high energy (VHE) photons (e.g., Albert et al., 2007; Aharonian et al., 2007) and have been proposed as efficient proton accelerators (starting with Berezinsky, 1977), capable of producing neutrinos at TeV–PeV and beyond.
 Their observed properties are strongly influenced by relativistic beaming effects that also introduce significant observational biases.
-For example, flux-density-limited samples of radio Very Long Baseline Interferometry (VLBI) selected AGN are dominated by distant beamed blazars with a median redshift of $z\sim 1$ (L.
-Y.
-Petrov & Y.
-Y.
-Kovalev, 2025).
+For example, flux-density-limited samples of radio Very Long Baseline Interferometry (VLBI) selected AGN are dominated by distant beamed blazars with a median redshift of $z\sim 1$ (Petrov & Kovalev, 2025).
 However, the role of beaming at very high energies ($\gtrsim 100$ GeV) is not firmly established.
 Indeed, while fast variability of VHE gamma-rays implies high values of the Doppler factor $\delta$, the lack of strong apparent superluminal motions in VLBI observed jets suggests the opposite.
-This is known as the ‘Doppler factor crisis’ (see, e.g., J.
-Albert et al., 2007; F.
-Aharonian et al., 2007; B.
-G.
-Piner & P.
-G.
-Edwards, 2018).
+This is known as the “Doppler factor crisis” (see, e.g., Albert et al., 2007; Aharonian et al., 2007; Piner & Edwards, 2018).
 Breakthroughs in this area have been rare over the past 35 years, with the Doppler factor crisis posing a significant barrier to progress.
-The neutrino emission from distant blazars is expected to be beamed, with proton acceleration occurring preferentially along the jet direction (e.g., IceCube Collaboration et al., 2018; A.
-V.
-Plavin et al., 2025).
-In contrast, weaker AGN may emit neutrinos (R.
-Abbasi et al., 2022) without any preferred direction.
+The neutrino emission from distant blazars is expected to be beamed, with proton acceleration occurring preferentially along the jet direction (e.g., IceCube Collaboration et al., 2018; Plavin et al., 2025).
+In contrast, weaker AGN may emit neutrinos (Abbasi et al., 2022) without any preferred direction.
 Figure 1: The IceCube skymap cutout surrounding the radio position of the PKS 1424+240 blazar.
-The color scale represents the local p-value from the 9-year maximum likelihood analysis performed by R.
-Abbasi et al.
+The color scale represents the local p-value from the 9-year maximum likelihood analysis performed by Abbasi et al.
 (2022).
-With a redshift $z=0.605$ (S.
-Paiano et al., 2017), the BL Lac object PKS 1424+240 (common name OQ 240 (catalog )) is the most distant blazar detected in quiescent VHE (e.g., J.
-Aleksić et al., 2014; P.
-Padovani et al., 2022).
-The all-sky analysis performed by R.
-Abbasi et al.
+With a redshift $z=0.605$ (Paiano et al., 2017), the BL Lac object PKS 1424+240 (common name OQ 240) is the most distant blazar detected in quiescent VHE (e.g., Aleksić et al., 2014; Padovani et al., 2022).
+The all-sky analysis performed by Abbasi et al.
 (2022) has identified it as the possible source of the second-highest neutrino excess in the predefined source list in the Northern sky (see Figure 1).
-The best-fit neutrino spectrum for PKS 1424+240 is soft with a spectral index of $\gamma=3.5$ (R.
-Abbasi et al., 2022).
+The best-fit neutrino spectrum for PKS 1424+240 is soft with a spectral index of $\gamma=3.5$ (Abbasi et al., 2022).
 The IceCube signal is dominated by neutrinos with relatively low, $\lesssim 10$ TeV, energies.
@@
 We use radio VLBI, which is the only technique to date that is able to directly probe the sites of neutrino and gamma-ray production, resolving structures within several parsecs from the central engine.
-For the redshift of PKS 1424+240, the projected angular scale of 1 mas corresponds to 6.70 pc, assuming the E.
-Komatsu et al.
+For the redshift of PKS 1424+240, the projected angular scale of 1 mas corresponds to 6.70 pc, assuming the Komatsu et al.
 (2009) cosmological parameters.
 A quick examination of PKS 1424+240 has revealed that its jet has an exceptionally large apparent opening angle $\varphi_{\mathrm{app}}$.
-A.
-B.
 Pushkarev et al.
-(2017) found $\varphi_{\mathrm{app}}=65\arcdeg$, among the most extreme values observed for blazars.
+(2017) found $\varphi_{\mathrm{app}}=65\degr$, among the most extreme values observed for blazars.
 Geometrically, this indicates a small viewing angle $\theta$ to the line of sight and, consequently, strong Doppler boosting.
@@
 The results of our serendipitous findings are presented in this paper.
-2 VLBA polarization imaging
-2.1 Long-term 15 GHz MOJAVE observations: image stacking
-Stokes I and polarization observation of PKS 1424+240 in the MOJAVE program [^1] https://www.cv.nrao.edu/MOJAVE/ (MOJAVE: Monitoring Of Jets in Active galactic nuclei with VLBA Experiments, see M.
-L.
-Lister et al., 2018, and references therein) began in May 2009 and we report results through January 2025.
-To increase the dynamic range of a restored map, a set of single-epoch images is combined and averaged, i.e., stacked, by aligning them using the VLBI core position derived from visibility model fitting (e.g., M.
-L.
-Lister et al., 2021).
+2 15 GHz VLBA polarization imaging
+Stokes I and polarization observation of PKS 1424+240 in the MOJAVE program [^1] https://www.cv.nrao.edu/MOJAVE/ (MOJAVE: Monitoring Of Jets in Active galactic nuclei with VLBA Experiments, see Lister et al., 2018, and references therein) began in May 2009 and we report results through January 2025.
+To increase the dynamic range of a restored map, a set of single-epoch images was combined and averaged, i.e., stacked, by aligning them using the VLBI core position derived from visibility model fitting (e.g., Lister et al., 2021).
 Stacking not only improves the sensitivity and dynamic range but also effectively reconstructs more complete source morphology, particularly in the low-brightness areas.
 It can also reveal the whole jet channel, not just regions of enhanced radiation at a given epoch.
-It is beneficial for studies of jets in Stokes I and polarized emission (e.g.
-A.
-B.
-Pushkarev et al., 2017, 2023; Y.
-Y.
-Kovalev et al., 2020; D.
-I.
-Zobnina et al., 2023).
-We made stacked maps of PKS 1424+240 in Stokes I, Q, and U (Figure 2), using 42 available epochs that span approximately 15 years, following a procedure described in A.
-B.
-Pushkarev et al.
+Image stacking is beneficial for studies of jets in Stokes I and polarized emission (e.g.
+Pushkarev et al., 2017, 2023; Kovalev et al., 2020; Zobnina et al., 2023).
+We made stacked maps of PKS 1424+240 in Stokes I, Q, and U (Figure 2), using 42 available epochs that span approximately 15 years, following a procedure described in Pushkarev et al.
 (2023).
-For deep CLEANing in Stokes I, Q, and U, we applied the algorithm with the residual entropy-based stopping criterion developed by D.
-C.
-Homan et al.
+For deep CLEANing in Stokes I, Q, and U, we applied the algorithm with the residual entropy-based stopping criterion developed by Homan et al.
 (2024).
-The linear polarization image has been corrected for Ricean bias (J.
-F.
-C.
-Wardle & P.
-P.
-Kronberg, 1974).
-We also applied a debiasing procedure to the CLEAN images to effectively suppress the sidelobe artifacts from deconvolution (Appendix B in A.
-B.
-Pushkarev et al., 2023).
+The linear polarization image has been corrected for Ricean bias (Wardle & Kronberg, 1974).
+We also applied a debiasing procedure to the CLEAN images to effectively suppress the sidelobe artifacts from deconvolution (Appendix B in Pushkarev et al., 2023).
 Figure 2: VLBA stacked image for the blazar PKS 1424+240 at 15 GHz.
@@
 A circular restoring beam is shown at the full width at half maximum (FWHM, 0.8 mas) level in the bottom left corner.
-FITS files of the stacked Stokes I, Q, and U images are published as electronic supplementary material.
-We also supply an animation of Stokes I cumulative stacking.
+FITS files of the stacked Stokes I, Q, and U images are available at the CDS.
+We also supply an animation of Stokes I cumulative stacking (available online).
 This debiasing procedure has significantly reduced imaging artifacts, and we argue that the weak structures observed around the core of PKS 1424+240 in the stacked images (Figure 2) represent genuine reconstructed jet emission.
@@
 The reduced levels of detectable 15 GHz Stokes I emission in the North-West direction next to the bright core component possibly results from residual effects of self-calibration which only included emission to the South-East in the self-calibration model in most individual epochs.
-In order to check for consistency of the stacked 15 GHz image, we have processed archival 1.6 GHz VLBA data for PKS 1424+240, experiment BS257A, epoch 9 May 2017.
-At ten times lower frequency, the VLBA is substantially more sensitive to distant optically thin jet emission and probes angular scales ten times larger than our MOJAVE stacked image.
-With a total on-source time of about 3 hr, these single epoch low-frequency observations reach a high dynamic range of $\sim 7,000:1$ and confirm the emission around the VLBI core, including the North-West direction (Figure 4).
-We note that our stacked image at 24 GHz also detects this emission (Figure 5).
-2.2 Rotation Measure at 15, 24, and 43 GHz
-Between August 2019 and June 2021, the MOJAVE monitoring conducted monthly 15 GHz, 24 GHz, and 43 GHz band polarimetric observations of 25 AGN (including PKS 1424+240), totaling 23 epochs, in order to measure the rotation measure (RM; full results will be presented by Livingston and the MOJAVE team, in prep.).
-In Figure 5 we show the mean RM of PKS 1424+240 across the two years of observations.
-This object is one of the lowest in total magnitude of observed RM for its core, with a median magnitude across all epochs of $302\pm 108\,\mathrm{rad\,m^{-2}}$, compared to a population median magnitude of $1026\pm 475\,\mathrm{rad\,m^{-2}}$.
-We also note that the median pixel-by-pixel standard deviation of RM for the core is $514\pm 57\,\mathrm{rad\,m^{-2}}$, making it larger than the median magnitude of RM for the core.
-Interestingly, across the two years, the core of PKS 1424+240 does not vary greatly in the magnitude of RM, sitting in the bottom $\mathrm{25^{th}}$ percentile of time variability of sources.
-However, the jet of PKS 1424+240 has a time variability of RM similar to other sources within the sample.
-We also detect frequency dependent depolarization, which is typical of most sources within our sample.
-This combined with the low magnitude RM suggests that the Faraday rotation that we do see is external to the jet.
-As RM data is not available for all 42 epochs used in Figure 2, and individual epochs do not have RM data for the full jet region, we cannot apply a precise correction for the Faraday effect on EVPA.
-The median magnitude of RM for PKS 1424+240 corresponds to a EVPA rotation of $\sim 6^{\circ}$, which is comparable to our EVPA errors from calibration and imaging.
-The low magnitude and variability of RM especially close to the core of PKS 1424+240 are indicative of either a weak magnetic field or low thermal electron density along the line-of-sight.
-We may be seeing the ‘sweeping away’ of the magneto-ionic medium, this would result in a reduction in the magnitude of RM.
-This effect would be more prominent within the center of the jet, while the edges would be less affected, which is supported by the magnitude and variability of RM being typical of other sources for the jet region of PKS 1424+240.
+In order to check for consistency the Figure 2 Stokes I image, we have processed archival 1.6 GHz VLBA data for PKS 1424+240, see subsection A.1.
 3 Looking into the jet cone
@@
 Furthermore, the stacked polarization image exhibits a remarkably uniform distribution of EVPAs; it presents a pattern of “diverging rays”, which start from the core.
+Rotation measure 15/24/43 GHz VLBA results are presented in subsection A.2.
 The low rotation measure and its minimal variation suggest that the jet is propagating toward the observer, reducing the density of the interstellar medium along its path.
@@
 We reconstruct the magnetic field structure, implementing a $90^{\circ}$ EVPA rotation across the entire map.
-This assumes that RM as well as relativistic effects do not affect the observed projected magnetic field direction significantly, see subsection 2.2 and M.
-Lyutikov et al.
+This assumes that RM as well as relativistic effects do not affect the observed projected magnetic field direction significantly, see subsection A.2 and Lyutikov et al.
 (2005).
@@
 Tolkien) we call ‘the Eye of Sauron’ due to its striking nature.
-Linear Integral Convolution (LIC; B.
-Cabral & L.
-C.
-Leedom, 1993) is applied to the polarized intensity to visualize the morphology of magnetic field lines by locally blurring intensity variations along the magnetic field direction, effectively tracing its structure.
+Linear Integral Convolution (LIC; Cabral & Leedom, 1993) is applied to the polarized intensity to visualize the morphology of magnetic field lines by locally blurring intensity variations along the magnetic field direction, effectively tracing its structure.
 The image reveals a significant toroidal magnetic field component of a current-carrying jet, which flows almost directly toward Earth.
-This is consistent with a widely discussed overall helical magnetic field structure (R.
-Blandford et al., 2019).
+This is consistent with a widely discussed overall helical magnetic field structure (Blandford et al., 2019).
 3.2 Basic jet parameters
-M.
-L.
 Lister et al.
@@
 Their apparent speed, measured in units of the speed of light, is $\beta_{\mathrm{app,1}}=2.83\pm 0.89$ and $\beta_{\mathrm{app,2}}=1.91\pm 0.18$.
-These moderate values are consistent with other measurements for the same source (e.g., P.
-Padovani et al., 2022; E.
-Kun & A.
-Medveczky, 2023).
-A median intrinsic full opening angle of the jet was estimated by A.
-B.
-Pushkarev et al.
+These moderate values are consistent with other measurements for the same source (e.g., Padovani et al., 2022; Kun & Medveczky, 2023).
+A median intrinsic full opening angle of the jet was estimated by Pushkarev et al.
 (2017) to be $\alpha_{\mathrm{int}}=1.2^{\circ}$ for Fermi-detected AGN in MOJAVE sample.
@@
 (in prep.).
-We also assume a relation between $\alpha_{\mathrm{int}}$ and $\Gamma$, as deduced observationally by A.
-B.
-Pushkarev et al.
-(2017) and predicted by hydrodynamical and magnetic acceleration models of relativistic jets (R.
-D.
-Blandford & A.
-Königl, 1979; S.
-S.
-Komissarov et al., 2007; E.
-Clausen-Brown et al., 2013).
+We also assume a relation between $\alpha_{\mathrm{int}}$ and $\Gamma$, as deduced observationally by Pushkarev et al.
+(2017) and predicted by hydrodynamical and magnetic acceleration models of relativistic jets (Blandford & Königl, 1979; Komissarov et al., 2007; Clausen-Brown et al., 2013).
 As a result, we find that the following set of parameters for a basic relativistic jet model are consistent with the above values and considerations: $\Gamma=16$ (plasma flow speed $\beta=0.998$), $\alpha_{\mathrm{int}}=1.2^{\circ}$, $\theta=0.3^{\circ}$, $\beta_{\mathrm{app}}=2.8$, and Doppler factor $\delta=32$.
-The Lorentz factor is within its typical values from MOJAVE population modeling (M.
-L.
-Lister et al., 2019), as expected from the chosen conservative parameter assumptions.
-The Doppler factor is three times higher than the median values for the MOJAVE sample (M.
-L.
-Lister et al., 2021; D.
-C.
-Homan et al., 2021), which agrees with recent finding of A.
-V.
-Plavin et al.
+The Lorentz factor is within its typical values from MOJAVE population modeling (Lister et al., 2019), as expected from the chosen conservative parameter assumptions.
+The Doppler factor is three times higher than the median values for the MOJAVE sample (Lister et al., 2021; Homan et al., 2021), which agrees with recent finding of Plavin et al.
 (2025) for neutrino-selected blazars.
@@
 Assuming a jet viewing angle in the range $0.1^{\circ}<\theta<1.0^{\circ}$ and using the measured $\beta_{\mathrm{app}}$ values, the Doppler factors lie within the range $1652>\delta>$ 16.
-Although the geometric probability of a near-zero viewing angle is low, the fraction of such jets in VLBI flux density-selected samples has been shown to be on the order of a few percent due to Doppler bias (M.
-L.
-Lister et al., 2019).
+Although the geometric probability of a near-zero viewing angle is low, the fraction of such jets in VLBI flux density-selected samples has been shown to be on the order of a few percent due to Doppler bias (Lister et al., 2019).
 Due to strong projection effects and synchrotron opacity, the jet is expected to become visible many parsecs downstream from its true base.
 As a result, its observed radio core appears less active and variable than typically expected for highly boosted blazars.
-Moreover, the core and jet emission are superimposed, resulting in a significant underestimation of the intrinsic core brightness temperature, which in turn leads to an incorrect estimation of jet parameters in D.
-C.
-Homan et al.
+Moreover, the core and jet emission are superimposed, resulting in a significant underestimation of the intrinsic core brightness temperature, which in turn leads to an incorrect estimation of jet parameters in Homan et al.
 (2021).
+A moderate flare occurred in the core of PKS 1424+240 during 2019–2020; the core flux density has reached 0.27 Jy, while the median value over all epochs is 0.18 Jy.
+It did not affect these estimates significantly.
 4 Emission of high energy photons and neutrinos: relativistic beaming
 […]
-We note that the explanation of the VHE gamma-ray flux within a leptohadronic model of M.
-Cerruti et al.
+We note that the explanation of the VHE gamma-ray flux within a leptohadronic model of Cerruti et al.
 (2017) requires the Doppler factor of $\sim 30$, consistent with our observations.
 The expected associated neutrino flux is of order of that inferred from IceCube observations.
-The neutrino production in the blazar radio core may proceed through the $p\gamma$ mechanism, adopted by A.
-V.
-Plavin et al.
-(2021); O.
-Kalashev et al.
+The neutrino production in the blazar radio core may proceed through the $p\gamma$ mechanism, adopted by Plavin et al.
+(2021) and Kalashev et al.
 (2023) to explain TeV to PeV neutrino production.
-Energetic protons, required for the neutrino and, in the hadronic scenarios, gamma-ray production, may be accelerated either in the vicinity of the central black hole (K.
-Ptitsyna & A.
-Neronov, 2016) or at the border surface between the fast spine and the slower sheath in the jet (e.g., F.
-Tavecchio & G.
-Ghisellini, 2015).
-In any case, the specific geometry of the jet in PKS 1424+240, which is closely aligned to the line-of-sight, provides for an additional Doppler enhancement of its multi-wavelength and multi-messenger fluxes.
-This boosting makes the object persistently bright, maintaining a high average flux — placing it among the top 1 % of gamma-ray sources (J.
-Ballet et al., 2023), and the brightest blazar in terms of high-energy neutrino emission (R.
-Abbasi et al., 2022).
+Energetic protons, required for the neutrino and, in the hadronic scenarios, gamma-ray production, may be accelerated either in the vicinity of the central black hole (Ptitsyna & Neronov, 2016) or at the border surface between the fast spine and the slower sheath in the jet (e.g., Tavecchio & Ghisellini, 2015).
+In any case, the specific geometry of the jet in PKS 1424+240, which is closely aligned to the line of sight, provides an additional Doppler enhancement of its multi-wavelength and multi-messenger fluxes.
+This boosting makes the object persistently bright, maintaining a high average flux — placing it among the top 1% of gamma-ray sources (Ballet et al., 2023), and the brightest blazar in terms of high-energy neutrino emission (Abbasi et al., 2022).
 Future observations of this and other VHE emitting blazars are needed to pave the way to a quantitative model of neutrino production in jets, and to better understand the role of beaming in the gamma-ray emission.
@@
 The mechanism responsible for accelerating massive particles to relativistic energies remains an open question in modern astrophysics, with powerful extragalactic jets being a key candidate.
-A longstanding discrepancy—known as the Doppler factor crisis—arises from the mismatch between low apparent plasma speeds observed with VLBI and the highly relativistic flows required to explain the emission.
+A longstanding discrepancy — known as the Doppler factor crisis — arises from the mismatch between low apparent plasma speeds observed with VLBI and the highly relativistic jet required to explain the emission.
 We use 16 years of 2 cm VLBI observations to uncover properties of the parsec-scale jet of the blazar PKS 1424+240, which is exceptional in its VHE gamma-ray and probable neutrino emissions.
-A significant toroidal magnetic-field is detected by a clear signature in linear polarization, revealing a current-carrying jet flowing almost directly toward us.
+A significant toroidal magnetic field is detected by a clear signature in linear polarization, revealing a current-carrying jet flowing almost directly toward us.
 We determine that the jet is observed within its cone, at a viewing angle below half a degree.
-This geometry maximizes relativistic boosting while minimizing apparent speed, thereby resolving the Doppler factor crisis and suggesting extreme relativistic beaming for this VHE photon and neutrino emitting blazar.
+This geometry maximizes relativistic boosting while minimizing apparent speed, thereby resolving the Doppler factor crisis and suggesting extreme relativistic beaming for this VHE photon and neutrino-emitting blazar.
 Our findings for PKS 1424+240 pave the way towards better understanding of high-energy photon and neutrino emission from blazar jets.
@@
 Moreover, such jets may represent a distinct class of AGN with both enhanced VHE gamma-ray and neutrino emission.
-From Monte Carlo simulations of the complete, VLBI-selected MOJAVE 1.5JyQC sample, we estimate that only a few percent of the jets are viewed within a degree to our line of sight (M.
-L.
-Lister et al., 2019).
+From Monte Carlo simulations of the complete, VLBI-selected MOJAVE 1.5JyQC sample, we estimate that only a few percent of the jets are viewed within a degree to our line of sight (Lister et al., 2019).
 The results of a full MOJAVE sample analysis will be presented separately by Pushkarev et al.
 (in prep.).
+6 Data availability
+Stacked Stokes I, Q, U 15 GHz FITS images (Figure 2) as well as Stokes I 1.6 GHz FITS image (Figure 4) are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
+Acknowledgements.
 We thank Andrei Lobanov, Yannis Liodakis, and Tigran Arshakian for productive discussions and comments on the manuscript.
@@
 AVP is a postdoctoral fellow at the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (grants 60477, 61479, 62286) and the Gordon and Betty Moore Foundation (grant GBMF-8273).
-The work of SVT and ABP is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541.
+The work of SVT is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541.
 The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations.
@@
 This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
-This research made use of the data from the MOJAVE database maintained by the MOJAVE team (M.
-L.
-Lister et al., 2018).
-Appendix A Supplementary Figures
+This research made use of the data from the MOJAVE database maintained by the MOJAVE team (Lister et al., 2018).
+Appendix A Supporting observational material
+A.1 VLBA data at 1.6 GHz
+We have processed 1.6 GHz VLBA observations of PKS 1424+240, epoch 9 May 2017, experiment code BS257A.
+At ten times lower frequency, the VLBA is substantially more sensitive to distant optically thin jet emission and probes angular scales ten times larger than our MOJAVE stacked image.
+With a total on-source time of about 3 hr, these single epoch low-frequency observations reach a high dynamic range of about 7,000:1 and confirm the emission around the VLBI core, including the North-West direction (Figure 4).
+We note that our stacked image at 24 GHz also detects this emission (Figure 5).
+A.2 Rotation Measure at 15, 24, and 43 GHz
+Between August 2019 and June 2021, the MOJAVE monitoring conducted monthly 15 GHz, 24 GHz, and 43 GHz band polarimetric observations of 25 AGN (including PKS 1424+240), totaling 23 epochs, in order to measure the rotation measure (RM; full results will be presented by Livingston and the MOJAVE team, in prep.).
+In Figure 5 we show the mean RM of PKS 1424+240 across the two years of observations.
+This object is one of the lowest in total magnitude of observed RM for its core, with a median magnitude across all epochs of $302\pm 108\,\mathrm{rad\,m^{-2}}$, compared to a population median magnitude of $1026\pm 475\,\mathrm{rad\,m^{-2}}$.
+We also note that the median pixel-by-pixel standard deviation of RM for the core is $514\pm 57\,\mathrm{rad\,m^{-2}}$, making it larger than the median magnitude of RM for the core.
+Interestingly, across the two years, the core of PKS 1424+240 does not vary greatly in the magnitude of RM, sitting in the bottom $\mathrm{25^{th}}$ percentile of time variability of sources.
+However, the jet of PKS 1424+240 has a time variability of RM similar to other sources within the sample.
+We also detect frequency dependent depolarization, which is typical of most sources within our sample.
+This combined with the low magnitude RM suggests that the Faraday rotation that we do see is external to the jet.
+As RM data is not available for all 42 epochs used in Figure 2 and individual epochs do not have RM data for the full jet region, we cannot apply a precise correction for the Faraday effect on EVPA.
+The median magnitude of RM for PKS 1424+240 corresponds to a EVPA rotation of $\sim 6^{\circ}$, which is comparable to our EVPA errors from calibration and imaging.
+The low magnitude and variability of RM especially close to the core of PKS 1424+240 are indicative of either a weak magnetic field or low thermal electron density along the line of sight.
+We may be seeing the ‘sweeping away’ of the magneto-ionic medium, this would result in a reduction in the magnitude of RM.
+This effect would be more prominent within the center of the jet, while the edges would be less affected, which is supported by the magnitude and variability of RM being typical of other sources for the jet region of PKS 1424+240.
 Figure 4: 1.6 GHz VLBA image of PKS 1424+240, epoch 9 May 2017.
 A circular restoring beam is shown in the left bottom corner at the FWHM level of 8.17 mas.
-It supports the viewing angle being less than a half-opening angle even at large (kiloparsec, deprojected) spatial scales probed by the low-frequency observations.
-Stokes I image FITS file is published electronically as a paper supplementary material.
+The image supports the viewing angle being less than a half-opening angle even at large (kiloparsec, deprojected) spatial scales probed by the low-frequency observations.
+Stokes I image FITS file is is available at the CDS.
 Figure 5: Stacked mean RM map of 15/24/43 GHz VLBA data for PKS 1424+240 in the frame of the observer across 23 epochs from 15 August 2019 to 25 June 2021 in color scale.